OpenStudy (cenaida):

Simplify (10x4y2) ÷ (5xy). A. 80 B. 5x3y C. 23yx3 D. 2x3y

OpenStudy (anonymous):

(10/5)*(x^4/x)*(y^2)/y

OpenStudy (anonymous):

\[(10x^4y^2) ÷ (5xy) = \frac {10x^4y^2} { 5xy }=\frac{10}{5}x^{4-1}y^{2-1}\] \[=2x^{3}y^{1}=2x^{3}y\] Thus option D. \[2x^3y\] is the correct Answer. @cenaida